OpenStudy (anonymous):

f(x)= x^3+6x^2+x^1/2 and g(x)= x^1/2. Find f(x)/g(x). A. x^5/2+6x^3/2-1 B. (x^5+6x^3+1)^1/2 C. x^5/2+6x^3/2+1 D. x^3/2+6x+x^1/4 C. x^3/2-6x+x^1/4

OpenStudy (anonymous):

Ich habe keine Ahnung , was du redest .

OpenStudy (anonymous):

What?

imqwerty (imqwerty):

c

OpenStudy (fibonaccichick666):

so you need to simplify this yea? \[f(x)= x^3+6x^2+x^{1/2}~~ and~~ g(x)= x^{1/2}.\\ \frac{f(x)}{g(x)}=\frac{x^3+6x^2+x^{1/2}}{x^{1/2}}\]